OpenStudy (anonymous):

An airplane's velocity is recorded at 5-min intervals during a 1-hour period with the following results, in miles per hour: 625; 640; 570; 580; 565; 550; 570; 600; 575; 595; 600; 575; 605; Use Simpson’s Rule to estimate the distance traveled during the hour. Answer in miles (using mi for miles) to the nearest whole mile.

OpenStudy (anonymous):

Do you remember Simpson's rule?

OpenStudy (anonymous):

It's a way to approximate integrals.

OpenStudy (anonymous):

Yes, but I keep getting 588.056 miles, but I know that this is wrong. I can't tell what I messed up on.

OpenStudy (anonymous):

I would suggest laying these out on a graph with a=550 to b=640; you then want to find the change in X of the interval: \[\Delta x =\frac{b-a}n\] Where n is the subintervals of equal length. Essentially, you have something like this: \[\int\limits_{550}^{640}f(x)dx= \frac{\Delta}3(f(x_0)+4f(x_1)+2f(x_2)+ \cdots + f(x_n))=S_n\] Now just plug the values in to the formula and you should have the solution easily.

OpenStudy (anonymous):

I'm guessing n=12?

OpenStudy (anonymous):

If I take all of the repeats out, I get 10 data values, so n=10?

OpenStudy (anonymous):

No. I erased the last comment because it was confusing. Sorry. You should have n=12 and 12 x-values (they correspond). However, you have 13 values.